# Sample Complexity Analysis of Multi-Target Detection via Markovian and Hard-Core Multi-Reference Alignment

### v2

@@ 6. The two-dimensional MTD model: Main results
Analogously to Section 3, we present the main results for the two–dimensional hard-core MTD model introduced in the previous section. In Section 6.1, we show that an MTD measurement can be partitioned into $L\times L$ patches, each of which has the form of a two–dimensional MRA observation (see Figure 2). Although these patches are not i.i.d., they are generated by an underlying random field that exhibits exponential mixing, governed by the distribution placement of the hard-core model (Assumption 5.3). In Section 6.2, we prove that this dependent MRA model converges to the classical i.i.d. MRA model; consequently, the sample complexity of the induced 2D hard-core model can be analyzed by comparison with the i.i.d. MRA framework.
@@
In contrast to the one–dimensional Markov case (cf. (16)), the distribution $\pi$ does not admit a simple closed form in 2D. Instead, it is constructed in the proof of Theorem 6.2 as follows:
@@
As in 1D, statement (ii) in Theorem 6.3 shows that empirical–average estimators (when applicable) need no subsampling and thus exploit all $M^{2}$ patches. Their MSE continues to scale as $a(\sigma)/M^{2}$ under exponential mixing, matching the i.i.d. rate up to constants. The same implication holds for the method of moments: empirical $n$th–order moments computed from the 2D MTD patches converge to their i.i.d. counterparts at rate $a(\sigma)/M^{2}$. If $n_{\min}$ is the minimal moment order needed for orbit recovery in the i.i.d. model, then in the low–SNR regime the required number of patches $N=M^{2}$ satisfies $N=\omega(\sigma^{2n_{\min}})$, identical scaling to the i.i.d. case.
@@ 7.1. Sample complexity for the MTD model
In this work, we establish a formal connection between the MTD model and the classical i.i.d. MRA problem. In the one-dimensional setting, the MTD model is cast as a Markovian variant of MRA, while in two dimensions, it is cast as a hard-core interaction model. We show that the MSE in the Markovian or hard-core setting can be upper bounded by that of the corresponding i.i.d. MRA model. The resulting i.i.d. MRA model falls within the standard MRA paradigm, where the sample complexity is known to depend on the minimal moment that uniquely determines the orbit of the underlying signal. While this minimal moment has been explicitly characterized for several standard MRA models, determining it for the i.i.d. model induced by our Markovian construction remains an open problem. A resolution of this question would enable a sharper and fully explicit upper bound on the sample complexity of the MTD model.
@@ 7.3. Connection to cryo-EM and cryo-ET
While extending our methodology to the cryo-EM setting poses challenges, primarily due to the complexities introduced by projection imaging, a more tractable and closely related direction arises in cryo-electron tomography (cryo-ET), particularly in the context of the sub-tomogram averaging problem. In this setting, projections are absent, and the observations more closely resemble the raw signal model studied in this work, making it a promising candidate for applying the techniques developed here.
@@ Appendix A Proofs
If $Y_{k}$ contains the remainder of a signal which started in the previous patch, let’s say at the $l$-th pixel for some $l\in\{1,\ldots,L-1\}$, then we set $g_{k}^{(1)}=l$, i.e. $g_{k}^{(1)}\cdot\tilde{X}$ is the signal $\tilde{X}$, shifted $l$ pixels to the right. In this way, $P_{0}\left[g_{k}^{(1)}\cdot\tilde{X}\right$] contains the remainder of the signal that started in the previous patch. If, on the contrary, $Z_{k}$ has no signal remaining from the previous patch, then we take $g_{k}^{(1)}=0$. In this case, $P_{0}\left[g_{k}^{(1)}\cdot\tilde{X}\right]=P_{0}\tilde{X}=0_{L}$ is the vector of zeros.
@@
Step 3: Next, we prove that the Markov Chain $\{\omega_{k}\}_{k\geq 0}$ defined in (38)–(39) has a positive absolute spectral gap. For this, it is enough to verify that the Markov Chain is indecomposable and aperiodic (see [34, Section 4]). These two properties hold since, in view of the transition matrix $P$ in (39), the state $\omega=L$ is accessible (with positive probability) from any other state, and from this state, the Markov Chain can transition (with positive probability) to any other state.
This argument proves the existence of a unique stationary distribution $\rho$ on $\{0,\ldots,L\}$, and that, if we call $\mu_{k}(\omega|\omega_{m})$, with $m,k\geq 0$ the probability distribution of $\omega_{m+k}$ conditionally on $\omega_{m}$, then the probability distribution $\mu_{k}(\cdot|\omega_{m})$ converges to $\rho$ as $k$ tends to infinity. Using a coupling argument (see for instance [34, Section 6]) and the uniform minorization condition from [34, Section 6.2], we can obtain explicitly the following convergence estimate:
@@
Using this convergence estimate for the Markov chain $\{\omega_{k}\}_{k\geq 1}$ and the choice of $g_{k}$ in (40) and (41), we obtain in Lemma A.1 the stationary distribution associated to $\{g_{k}\}_{k\geq 1}$ and deduce the statement (iii) of the Theorem. $\square$
@@
introduced in Step 3 of the proof of Theorem 3.4. Hence, combining this with the definition of $\tilde{\rho}$ in (44) and (42), we obtain the following convergence estimate for the Markov chain $(\omega_{k},\omega_{k-1})$:
@@
    Since the Gibbs measure in $\mathbb{Z}^{2}$ is shift invariance, the restriction to any rectangle $V_{k}\subset V_{M}$ is the same for every $k\in\mathcal{I}_{M}$. Then, we use this restriction of the Gibbs measure to define, as in Step (ii), a probability distribution $\pi$ over the group $G$. Using the strong spatial mixing from the previous step, we prove that the random field $\left\{g_{k}\right\}_{k\in\mathcal{I}_{M}}$ and the probability distribution $\pi$ satisfy the exponential mixing property in Definition 5.5.
@@
Step 2: Let us recall the graph $\mathcal{G}_{M}=(V_{M},E_{M})$, with vertices $V_{M}\subset\mathbb{Z}^{2}$ given by (24) and edges $E_{M}\subset V_{M}\times V_{M}$ defined in (25). Let us denote by $\Theta$ the random variable associated to the hard-core model on the graph $\mathcal{G}_{M}$ with activity parameter $\lambda\in(0,1)$; see Definition 5.2. The hard-core model $\Theta$ is a random variable that takes values in the independent subgraphs of the graph $\mathcal{G}_{M}$, i.e. subsets of $V_{M}$ which contain no adjacent vertices. From now on, we represent subsets of $V_{M}$ by means of a boolean matrix of the form $\eta=\{\eta_{v}\ :\ v\in V_{M}\}\in\{0,1\}^{V_{M}}$. We recall that, for any $\eta\in\{0,1\}^{V_{M}}$, we say that $\eta$ is independent if and only if
@@
$\mathbb{P}(\Theta=\eta)=\begin{cases}C^{-1}\lambda^{|\eta|}&\text{if $\eta$ is independent},\\ 0&\text{else,}\end{cases}$
@@
$\mathbb{P}[B]=\mathbb{P}\left[\overline{\Theta}_{V_{S}}=\tau\mid\overline{\Theta}_{V^{c}_{M}}=0\right]=\dfrac{\mathbb{P}\left[(\overline{\Theta}_{V_{S}},\overline{\Theta}_{V_{M}^{c}})=(\tau,0)\right]}{\mathbb{P}\left[\overline{\Theta}_{V_{M}^{c}}=0\right]}$
@@
for some projection operator $P$ and some group elements $\{g_{k}\}_{k\in\mathcal{I}_{M}}$ in $G$, acting on the target signal $\overline{X}$, and $\varepsilon_{k}$ being i.i.d. Gaussian noise with variance $\sigma^{2}>0$. We also recall, from the statement of the theorem, the definition of the i.i.d. measurements $\{Y_{k}\}_{k\in\mathcal{I}_{M}}$, which have the same form of (55), but with group elements $\{\tilde{g}_{k}\}_{k\in\mathcal{I}_{M}}$ being i.i.d. from a suitable probability distribution $\pi$ over $G$.
Part (i): Let $c>0$ be a constant to be chosen later. For any $M\in\mathbb{N}$, choose $m=\lfloor c\log(M)\rfloor$ and $M^{\prime}=\lfloor M/m\rfloor$. We define the index set
@@
We claim that there exists $c_{1}>0$, independent of $M^{\prime}$, such that the $\ell_{1}$ distance between $\tilde{\pi}_{M}$ and $\pi_{M}$ satisfies

### v3

@@ 6. The two-dimensional MTD model: Main results
Analogously to Section 3, we present the main results for the two-dimensional hard-core MTD model introduced in the previous section. In Section 6.1, we show that an MTD measurement can be partitioned into $L\times L$ patches, each of which has the form of a two-dimensional MRA observation (see Figure 3). Although these patches are not i.i.d., they are generated by an underlying random field that exhibits exponential mixing, governed by the distribution placement of the hard-core model (Assumption 5.3). In Section 6.2, we prove that this dependent MRA model converges to the classical i.i.d. MRA model; consequently, the sample complexity of the induced 2D hard-core model can be analyzed by comparison with the i.i.d. MRA framework.
@@
In contrast to the one-dimensional Markov case (cf. (17)), the distribution $\pi$ does not admit a simple closed form in 2D. Instead, it is constructed in the proof of Theorem 6.2 as follows:
@@
As a noticeable difference with respect to the 1D case, in the 2D case we require the activity parameter $\lambda$ to be smaller than a constant $\lambda_{0}$, which depends on the degree of the graph defining the hard core model. In [42, Corollary 2.6]), the threshold $\lambda_{0}$ is given explicitly as $\lambda_{0}=\frac{b^{b}}{(b-1)^{b+1}}$, where $b+1$ is the degree of the graph (the maximum number of neighbors of each node), and one can deduce the simpler lower estimate $\lambda_{0}\geq\frac{1}{b-1}$. In view of Definition 5.2, the degree of the graph is $(2L-1)^{2}-1$, hence
$\lambda_{0}\geq\dfrac{1}{(2L-1)^{2}-2}.$
The result in [42, Corollary 2.6]) applies to general graphs, and we are not able to determine whether this condition is sharp for the specific type of graph that we are considering here. The necessity of the smallness condition on the activity parameter $\lambda$ to ensure exponential mixing is left as an open problem.
@@
As in 1D, statement (ii) in Theorem 6.3 shows that empirical–average estimators (when applicable) need no subsampling and thus exploit all $M^{2}$ patches. Their MSE continues to scale as $a(\sigma)/M^{2}$ under exponential mixing, matching the i.i.d. rate up to constants. The same implication holds for the method of moments: empirical $n$th–order moments computed from the 2D MTD patches converge to their i.i.d. counterparts at rate $a(\sigma)/M^{2}$. If $n_{\min}$ is the minimal moment order needed for orbit recovery in the i.i.d. model, then in the low-SNR regime the required number of patches $N=M^{2}$ satisfies $N=\omega(\sigma^{2n_{\min}})$, identical scaling to the i.i.d. case.
7. Numerical experiments
In Section 4, we established a precise statistical connection between the MTD model, its Markovian MRA formulation, and the corresponding induced i.i.d. MRA model. In particular, Theorem 4.2 and Corollary 4.3 show that, for empirical-average estimators such as empirical moments, the sample complexity in the MTD/Markovian MRA setting matches that of the induced i.i.d. model up to constant factors. Moreover, this connection relies on the convergence of the latent Markov chain to its stationary distribution, which governs the induced i.i.d. model and underlies the moment-based transfer of sample-complexity bounds.
The numerical experiments in this section serve three purposes. First, we empirically validate the stationary distribution and quantify the convergence of the latent Markov chain to stationarity. Second, we provide evidence that moments up to order $3$ suffice for recovery in the induced i.i.d. model, whereas second-order information alone is insufficient in the regimes tested here, thereby supporting the conclusion that $n_{\min}=3$. Third, we compare the observable moments of the Markovian and induced i.i.d. models under additive noise, illustrating the predicted scaling laws.
7.1. Convergence to the stationary distribution
We next validate the explicit stationary distribution $\pi$ in (17) empirically, and quantify how quickly the Markov chain $\{g_{k}\}_{k=1}^{M}$ approaches stationarity as the number of patches increases. For an observation composed of $M$ patches, we form the empirical distribution
(38) $\widehat{\pi}_{M}(g)\triangleq\frac{1}{M}\sum_{k=1}^{M}\mathbf{1}\{g_{k}=g\},\qquad g\in\{0,\ldots,L-1\}\times\{0,\ldots,L\},$
where, in the simulations, the latent states $g_{k}=(g_{k}^{(1)},g_{k}^{(2)})$ are computed from the ground-truth signal start times $\{t_{i}\}$ used to generate the MTD measurement (see the construction in Theorem 3.4 and the discussion in Section 3.2). We quantify the discrepancy between $\widehat{\pi}_{M}$ and the theoretical stationary distribution $\pi$ using the total variation distance
(39) $\mathrm{TV}\!\left(\widehat{\pi}_{M},\pi\right)\;\triangleq\;\frac{1}{2}\sum_{g}\left|\widehat{\pi}_{M}(g)-\pi(g)\right|.$
To connect convergence of $\widehat{\pi}_{M}$ to the method-of-moments viewpoint (Section 2.1), we also track convergence at the level of induced noiseless moments. Specifically, we form the empirical moment estimator $\widehat{T}_{\mathrm{MTD}}^{(2)}\triangleq\frac{1}{M}\sum_{k=1}^{M}\big(Z_{k}\big)^{\otimes 2}$, as defined in (19), for the observations $\{Z_{k}\}_{k=1}^{M}$ defined in (2) (without the noise term), and compare them to the corresponding noiseless population moments $T^{(2)}_{\overline{X},\pi}$ defined in (5), which are evaluated exactly by summation over $\operatorname{supp}(\pi)$. For each $(\lambda,M)$, we repeat the simulation over multiple independent Monte-Carlo trials and report the empirical means of $\mathrm{TV}(\widehat{\pi}_{M},\pi)$ and of the MSE of $\mathbb{E}\big\|\widehat{T}_{\mathrm{MTD}}^{(2)}-T^{(2)}_{\overline{X},\pi}\big\|_{F}^{2}$.
Figure 4 summarizes the results. Panel (a) plots the Monte-Carlo average of $\mathrm{TV}(\widehat{\pi}_{M},\pi)$ as a function of $M$, for several values of $\lambda$. Across all tested $\lambda$ values, the curves exhibit an approximately linear decay with slope close to 1/2-1/2 on log–log axes, indicating the scaling
(40) $\mathbb{E}\,\mathrm{TV}\!\left(\widehat{\pi}_{M},\pi\right)\;\asymp\;\frac{C(\lambda,L)}{\sqrt{M}},$
for a constant $C(\lambda,L)$ depending on the mixing properties of the chain (and therefore on $\lambda$ and $L$). Panel (b) report the corresponding moment estimation error and exhibit a $M^{-1}$ scaling for the moment MSEs:
(41) $\mathbb{E}\Big\|\widehat{T}_{\mathrm{MTD}}^{(2)}-T^{(2)}_{\overline{X},\pi}\Big\|_{F}^{2}\;\asymp\;\frac{C_{2}(\lambda,L,X)}{M},$
with constants depending on $(\lambda,L)$ (and on the fixed signal $X$). In addition, it can be seen that a smaller $\lambda$ enhances the convergence to the stationary distribution, as discussed in Section 3.3. These empirical observations are consistent with the exponential mixing bound in (16): as $M$ grows, the effective number of approximately independent samples increases linearly with $M$, yielding $M^{-1/2}$ fluctuations for TV distance and the corresponding $M^{-1}$ scaling for MSE of the moment estimation.
Figure 4. Empirical convergence to stationarity and induced moment scaling. (a) For each $\lambda$, we simulate the Markov chain $\{g_{k}\}$ induced by the 1D MTD patching procedure (Theorem 3.4), estimate its stationary distribution by $\widehat{\pi}_{M}$ in (38), and report the Monte-Carlo mean of the total variation distance $\mathrm{TV}(\widehat{\pi}_{M},\pi)$ in (39). (b) Monte-Carlo means of $\mathbb{E}\|\widehat{T}^{(2)}-T^{(2)}_{\overline{X},\pi}\|_{F}^{2}$, where $T^{(2)}_{\overline{X},\pi}$ denotes the noiseless population moments and $\widehat{T}_{\mathrm{MTD}}^{(2)}$ the corresponding empirical noiseless estimators. The dashed reference lines indicate slopes 1/2-1/2 in (a) and 1-1 in (b), consistent with the observed $M^{-1/2}$ scaling for TV and $M^{-1}$ scaling for moment MSE.
7.2. Empirical identification of the minimal moment order in the induced i.i.d. MRA model
This subsection provides empirical evidence regarding the moment cut-off order $n_{\min}$ of the induced i.i.d. MRA model associated with the MTD construction. In particular, our goal is to examine whether recovery is already possible from moments up to order $3$, or whether higher-order moment information may be necessary.
[…]
In our experiments, we generate a ground-truth signal $X_{\star}$, with i.i.d. Gaussian entries, compute the target second-order and third-order moments $\big(T^{(2)}_{\overline{X}_{\star},\pi},\,T^{(3)}_{\overline{X}_{\star},\pi}\big)$ via (5), and then attempt recovery by moment matching using either $T^{(2)}_{\overline{X}_{\star},\pi}$ alone or the pair $\big(T^{(2)}_{\overline{X}_{\star},\pi},\,T^{(3)}_{\overline{X}_{\star},\pi}\big)$. More precisely, we estimate $X$ by solving the nonconvex moment-matching problem
(42) $\widehat{X}\;\in\;\arg\min_{X\in\mathbb{R}^{L}}\Bigg\{\frac{\big\|T^{(2)}_{\overline{X},\pi}-T^{(2)}_{\overline{X}_{\star},\pi}\big\|_{F}^{2}}{\big\|T^{(2)}_{\overline{X}_{\star},\pi}\big\|_{F}^{2}}\;+\;w_{3}\,\frac{\big\|T^{(3)}_{\overline{X},\pi}-T^{(3)}_{\overline{X}_{\star},\pi}\big\|_{F}^{2}}{\big\|T^{(3)}_{\overline{X}_{\star},\pi}\big\|_{F}^{2}}\Bigg\},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm. We consider both the case $w_{3}>0$ (specifically, $w_{3}=3$) and the second-moment-only variant $w_{3}=0$.
We minimize (42) using fminunc of Matlab with multiple random restarts. Each restart is initialized with $X^{(0)}\sim\mathcal{N}(0,I_{L})$ and then normalized to satisfy $\|X^{(0)}\|_{2}=1$, thereby fixing the global scale. We run $n_{\mathrm{restarts}}$ independent initializations and report the solution attaining the smallest objective value. Since the optimization problem is non-convex, this procedure does not provide a theoretical guarantee of convergence to the global optimum, even with multiple restarts. Nevertheless, similar non-convex moment-matching formulations, solved via local optimization with random initialization, are commonly used in closely related recovery problems; see, for example, [19, 15]. The reference implementation is provided in https://github.com/AmnonBa/sample-complexity-mkv-mtd.
The results are shown in Figure 5. When matching moments up to order $3$ with $w_{3}=3$, the optimizer consistently returns an estimate $\widehat{X}$ that matches the target moments to numerical precision and achieves a small recovery error. By contrast, when matching only the second moment ($w_{3}=0$), the optimizer typically converges to alternative solutions that fit $T^{(2)}_{\overline{X}_{\star},\pi}$ well but still deviate substantially from $X_{\star}$, consistent with non-identifiability from second-order information alone in this induced model. We observe the same qualitative behavior across a range of density parameters $\lambda$ and for multiple signal lengths $L$, using randomly generated ground-truth signals. This indicates that the phenomenon is robust and not tied to a specific choice of $\lambda$ or dimension. Overall, these experiments provide empirical evidence that moments up to order $3$ suffice for recovery in the induced i.i.d. model in the regimes tested here, thereby supporting the conclusion that $n_{\min}=3$.
Based on these empirical observations, we make the following conjecture.
Conjecture 7.1 (Sample complexity of the induced i.i.d. MRA model).
Consider the induced i.i.d. MRA model associated with the one-dimensional MTD construction, under the assumptions of Section 3. Then, for a generic signal $X_{\star}$, the moment cut-off order is $n_{\min}=3$. Consequently, in the low-SNR regime, recovery of $X_{\star}$ up to the natural inherent group action requires sample complexity $M=\omega(\sigma^{6})$.
Figure 5. Recovery from second- and third-order moments versus second-order moments only. Noise-free population experiment in the induced i.i.d. MRA model. The black curve shows the ground-truth signal $X_{\star}$, and the red dashed curve shows the recovered estimate $\widehat{X}$ after optimal cyclic alignment. (a) Using both $\big(T^{(2)}_{\overline{X}_{,}\pi},T^{(3)}_{\overline{X},\pi}\big)$ yields accurate recovery. (b) Using $T^{(2)}_{\overline{X},\pi}$ alone leads to an ambiguous reconstruction that can match $T^{(2)}_{\overline{X},\pi}$ well but deviates substantially from $X_{\star}$, illustrating the need for third-order information in this induced model. In this simulation, we use $\lambda=0.3$, although similar behavior is observed for other values of $\lambda$.
7.3. Empirical Markovian-i.i.d. moment discrepancy under additive noise
This subsection empirically illustrates the connection established in Section 4 by quantifying, at the level of observable moments, how close the dependent MTD patch process (equivalently, the Markovian induced MRA model of Theorem 3.4) is to the induced i.i.d. MRA model with stationary distribution $\pi$. Whereas Theorem 4.2 guarantees i.i.d.-like variance scaling for empirical averages under Markovian sampling (up to constants determined by mixing), here we explicitly measure the finite-$M$ discrepancy between the two models, including its dependence on the noise level $\sigma$.
7.3.1. Models and moment estimators.
Fix $L$ and $\lambda$, and let $\pi$ be the stationary distribution in (17). We generate $M$ consecutive MTD patches $\{Z_{k}\}_{k=1}^{M}$ under Assumption 3.1 with additive noise $\varepsilon_{k}\sim\mathcal{N}(0,\sigma^{2}I_{L})$, and compute empirical moments $\widehat{T}^{(n)}_{\mathsf{MTD}}(M)$ for $n\in\{2,3\}$, as defined in (19). We then compare these to the population moments of the induced i.i.d. MRA model with $(g_{k}^{(1)},g_{k}^{(2)})\sim\pi$,
$T^{(n)}_{\mathsf{MRA}}(\sigma)\;\triangleq\;\mathbb{E}\!\left[Y^{\otimes n}\right],\qquad n\in\{2,3\},$
which we evaluate in the population limit by exact summation over $\operatorname{supp}(\pi)$ together with the standard Gaussian corrections for additive noise. We report the MSE between these moments by
(43) $\mathcal{E}_{n}(M,\sigma)\triangleq\mathbb{E}\!\left[\big\|\widehat{T}^{(n)}_{\mathsf{MTD}}(M)-T^{(n)}_{\mathsf{MRA}}(\sigma)\big\|_{F}^{2}\right],\qquad n\in\{2,3\},$
approximated via Monte-Carlo averaging over independent trials.
7.3.2. Observed scaling laws.
Figure 6 reports $\mathcal{E}_{2}(M,\sigma)$ and $\mathcal{E}_{3}(M,\sigma)$ as functions of $M$ for several noise levels. Three expected phenomena are visible. First, for fixed $\sigma$, the curves decay with slope close to 1-1 on log-log axes, consistent with the $1/M$ variance scaling for empirical averages, in agreement with Theorem 4.2(ii). Second, the separation across noise levels matches the standard noise-dominated moment-estimation scaling,
(44) $\mathcal{E}_{2}(M,\sigma)\;\propto\;\frac{\sigma^{4}}{M},\qquad\mathcal{E}_{3}(M,\sigma)\;\propto\;\frac{\sigma^{6}}{M},$
as the noise increases, as discussed in Section 2.1. Third, at very low noise the curves exhibit a nonzero floor attributable to finite-$M$ convergence of the Markov chain $\{g_{k}\}$ to stationarity: since $T^{(n)}_{\mathsf{MRA}}(\sigma)$ is defined under the stationary law $\pi$, any residual deviation of the state distribution from $\pi$ induces a bias in the Markovian moments, consistent with the total-variation convergence behavior in Figure 4(a).
Figure 6. Markovian-i.i.d. cross-model moment discrepancy under additive noise. We compare the empirical moments computed from $M$ consecutive MTD patches (Markovian induced MRA), $\widehat{T}^{(n)}_{\mathsf{MTD}}(M)=\frac{1}{M}\sum_{k=1}^{M}(Z_{k})^{\otimes n}$, to the population moments of the induced i.i.d. model under the stationary distribution $\pi$, $T^{(n)}_{\mathsf{MRA}}(\sigma)=\mathbb{E}[Y^{\otimes n}]$, and plot the Monte-Carlo estimates of $\mathcal{E}_{n}(M,\sigma)=\mathbb{E}\|\widehat{T}^{(n)}_{\mathsf{MTD}}(M)-T^{(n)}_{\mathsf{MRA}}(\sigma)\|_{F}^{2}$ for $n=2$ (panel a) and $n=3$ (panel b), across several noise levels $\sigma$. The dashed lines indicate the empirical $M^{-1}$ decay (variance of empirical averages), while the vertical separation between curves is consistent with the noise-dominated scaling $\mathcal{E}_{2}\propto\sigma^{4}/M$ and $\mathcal{E}_{3}\propto\sigma^{6}/M$. At very low noise, the curves exhibit a nonzero floor attributable to finite-$M$ convergence of the Markov chain state distribution to $\pi$ (cf. Figure 4(a)).
@@ 8. Discussion and outlook
8.1. Comparison with maximum-likelihood approaches
A natural alternative under the fully specified MTD model is maximum-likelihood estimation based on the marginal likelihood, obtained by integrating out the latent particle locations. While statistically appealing, this formulation is computationally challenging in our setting: even after marginalization, evaluating the likelihood requires summing over a combinatorially large collection of admissible latent configurations, which grows rapidly with the problem size. Existing related approaches therefore rely on approximate expectation-maximization / approximate marginal-likelihood formulations on simplified patch-level models rather than on exact maximum-likelihood [31, 32]. Such approximations can be effective computationally, but they still lead to nonconvex optimization problems and presently lack the kind of explicit low-SNR sample-complexity guarantees derived here. By contrast, our approach avoids direct likelihood optimization by reducing the problem to an induced MRA model, thereby enabling analysis through low-order moments and yielding explicit statistical guarantees.
@@ 8.2. Sample complexity for the MTD model
In this work, we establish a formal connection between the MTD model and the classical i.i.d. MRA problem. In the one-dimensional setting, the MTD model is cast as a Markovian variant of MRA, while in two dimensions, it is cast as a hard-core interaction model. We show that the MSE in the Markovian or hard-core setting can be upper bounded by that of the corresponding i.i.d. MRA model. The resulting i.i.d. MRA model falls within the standard MRA paradigm, where the sample complexity is known to depend on the minimal moment that uniquely determines the orbit of the underlying signal. While this minimal moment has been explicitly characterized for several standard MRA models, determining it for the i.i.d. model induced by our Markovian construction remains an open problem. In Section 7, however, we address this question empirically and provide evidence that moments up to order three suffice for recovery in the induced i.i.d. MRA model, resulting a sample complexity of $M=\omega(\sigma^{6})$. A rigorous characterization of the moment cut-off for this induced model remains an interesting direction for future work, and would lead to a sharper and fully explicit sample-complexity bound for the MTD model.
@@ 8.4. Connection to cryo-EM and cryo-ET
While extending our methodology to the cryo-EM setting poses challenges, primarily due to the complexities introduced by projection imaging, a more tractable and closely related direction arises in cryo-electron tomography (cryo-ET), particularly in the context of the sub-tomogram averaging problem. In this setting, tomographic projections are absent, and the observations more closely resemble the raw signal model studied in this work, making it a promising candidate for applying the techniques developed here.
@@ Appendix A Proofs
If $Z_{k}$ contains the remainder of a signal which started in the previous patch, let’s say at the $l$-th pixel for some $l\in\{1,\ldots,L-1\}$, then we set $g_{k}^{(1)}=l$, i.e. $g_{k}^{(1)}\cdot\tilde{X}$ is the signal $\tilde{X}$, shifted $l$ pixels to the right. In this way, $P_{0}\left[g_{k}^{(1)}\cdot\tilde{X}\right$] contains the remainder of the signal that started in the previous patch. If, on the contrary, $Z_{k}$ has no signal remaining from the previous patch, then we take $g_{k}^{(1)}=0$. In this case, $P_{0}\left[g_{k}^{(1)}\cdot\tilde{X}\right]=P_{0}\tilde{X}=0_{L}$ is the vector of zeros.
@@
Step 3: Next, we prove that the Markov Chain $\{\omega_{k}\}_{k\geq 0}$ defined in (46)–(47) has a positive absolute spectral gap. For this, it is enough to verify that the Markov Chain is indecomposable and aperiodic (see [37, Section 4]). These two properties hold since, in view of the transition matrix $P$ in (47), the state $\omega=L$ is accessible (with positive probability) from any other state, and from this state, the Markov Chain can transition (with positive probability) to any other state.
This argument proves the existence of a unique stationary distribution $\rho$ on $\{0,\ldots,L\}$, and that, if we call $\mu_{k}(\omega|\omega_{m})$, with $m,k\geq 0$ the probability distribution of $\omega_{m+k}$ conditionally on $\omega_{m}$, then the probability distribution $\mu_{k}(\cdot|\omega_{m})$ converges to $\rho$ as $k$ tends to infinity. Using a coupling argument (see for instance [37, Section 6]) and the uniform minorization condition from [37, Section 6.2], we can obtain explicitly the following convergence estimate:
@@
Using this convergence estimate for the Markov chain $\{\omega_{k}\}_{k\geq 0}$, we deduce a similar convergence result for the Markov chain $\{(\omega_{k},\omega_{k-1})\}_{k\geq 1}\subset\{0,\ldots,L\}^{2}$, with a stationary distribution denoted by $\tilde{\rho}$. Using the construction of $g_{k}$ in terms of $(\omega_{k},\omega_{k-1})$ given in (48) and (49), we obtain the convergence estimate in Theorem 3.4(iii) for the Markov chain $\{g_{k}\}_{k\geq 1}\subset G$. See Lemma A.1 for the explicit computations. The stationary distribution of the Markov chain $\{g_{k}\}_{k\geq 1}$ is denoted by $\pi$, and is simply the push-forward measure of $\tilde{\rho}$ by the map in $\{0,\ldots,L\}^{2}\to G$ defined in (49). $\square$
@@
introduced in Step 3 of the proof of Theorem 3.4. Hence, combining this with the definition of $\tilde{\rho}$ in (52), and the convergence estimate in (50), we obtain the following convergence estimate for the Markov chain $(\omega_{k},\omega_{k-1})$:
@@
    Since the Gibbs measure in $\mathbb{Z}^{2}$ is shift invariant, the restriction to any rectangle $V_{k}\subset V_{M}$ is the same for every $k\in\mathcal{I}_{M}$. Then, we use this restriction of the Gibbs measure to define, as in Step (ii), a probability distribution $\pi$ over the group $G$. Using the strong spatial mixing from the previous step, we prove that the random field $\left\{g_{k}\right\}_{k\in\mathcal{I}_{M}}$ and the probability distribution $\pi$ satisfy the exponential mixing property in Definition 5.5.
@@
Step 2: Let us recall the graph $\mathcal{G}_{M}=(V_{M},E_{M})$, with vertices $V_{M}\subset\mathbb{Z}^{2}$ given by (25) and edges $E_{M}\subset V_{M}\times V_{M}$ defined in (26). Let us denote by $\Theta$ the random variable associated to the hard-core model on the graph $\mathcal{G}_{M}$ with activity parameter $\lambda\in(0,1)$; see Definition 5.2. The hard-core model $\Theta$ is a random variable that takes values in the admissible subgraphs of the graph $\mathcal{G}_{M}$, i.e. subsets of $V_{M}$ which contain no adjacent vertices. From now on, we represent subsets of $V_{M}$ by means of a boolean matrix of the form $\eta=\{\eta_{v}\ :\ v\in V_{M}\}\in\{0,1\}^{V_{M}}$. We recall that, for any $\eta\in\{0,1\}^{V_{M}}$, we say that $\eta$ is admissible if and only if
@@
$\mathbb{P}(\Theta=\eta)=\begin{cases}C^{-1}\lambda^{|\eta|}&\text{if $\eta$ is admissible},\\ 0&\text{else,}\end{cases}$
@@
$\mathbb{P}[B]=\mathbb{P}\left[\overline{\Theta}_{V_{S}}=\tau\mid\overline{\Theta}_{V^{c}_{M}}=0\right]=\dfrac{\mathbb{P}\left[(\overline{\Theta}_{V_{S}},\overline{\Theta}_{V_{M}^{c}})=(\tau,0)\right]}{\mathbb{P}\left[\overline{\Theta}_{V_{M}^{c}}=0\right]}.$
@@
for some projection operator $P$ and some group elements $\{g_{k}\}_{k\in\mathcal{I}_{M}}$ in $G$, acting on the target signal $\overline{X}$, and $\varepsilon_{k}$ being i.i.d. Gaussian noise with variance $\sigma^{2}>0$. We also recall, from the statement of the theorem, the definition of the i.i.d. measurements $\{Y_{k}\}_{k\in\mathcal{I}_{M}}$, which have the same form as (63), but with group elements $\{\tilde{g}_{k}\}_{k\in\mathcal{I}_{M}}$ being i.i.d. from a suitable probability distribution $\pi$ over $G$.
The proof of the Theorems is naturally divided in two parts: part (i), which addresses the case of general estimators, considering an $m$-separated subsampling of the patches; and part (ii), which deals with estimators based on an empirical average, and do not require subsampling.
Part (i): We carry the proof in three steps. First, we unify the notation concerning the patch subsampling in the 1D and the 2D settings. Then, we prove an $\ell^{1}$-estimate for the probability distribution of the group elements associated the subsampled patches, which is a direct consequence of the mixing properties proved in Theorems 3.4 and 6.2. Finally, we combine the $\ell^{1}$-estimate with the law of total expectation to conclude the proof.
Step 1: Preparations. Let $c>0$ be a constant to be chosen later. For any $M\in\mathbb{N}$, choose $m=\lfloor c\log(M)\rfloor$ and $M^{\prime}=\lfloor M/m\rfloor$. We define the index set
@@
Step 2: Mixing property and $\ell^{1}$-estimate. We claim that there exists $c_{1}>0$, independent of $M^{\prime}$, such that the $\ell_{1}$ distance between $\tilde{\pi}_{M}$ and $\pi_{M}$ satisfies